6)# What is the limiting reactant for the following reaction given we have 3.4 moles of...

6)#
What is the limiting reactant for the following reaction given we have 3.4 moles of Ca(NO3)2 and 2.4 moles of Li3PO4?
Reaction: 3Ca(NO3)2 + 2Li3PO4 ---> 6LiNO3 + Ca3(PO4)2
Answer (show work please?)

Ca3(PO4)2

Li3PO4

Ca(NO3)2

LiNO3

not enough information
Answer #1

: 3Ca(NO3)2 + 2Li3PO4 ---> 6LiNO3 + Ca3(PO4)2

according to balanced reaction

3 moles Ca(NO3)2 reacts with 2 moles Li3PO4

3.4 moles Ca(NO3)2 reacts with 3.4 x 2 / 3 = 2.3 moles Li3PO4

but we have 2.4 moles of Li3PO4 . means exess Li3PO4 present

so exess reagent = Li3PO4

limiting reagent = Ca(NO3)2

answer = Ca(NO3)2